# Probing scalar non-standard interaction of supernova neutrinos in next-generation neutrino experiments

### v1

@@ 3 Neutrino flavor evolution in core-collapse supernova in the presence of flavor conserving SNSI
We now investigate the impact of scalar-mediated non-standard interactions on neutrino flavor evolution in the supernova environment. Specifically, we focus on flavor-diagonal SNSI parameters: $\eta_{ee}$, $\eta_{\mu\mu}$, and $\eta_{\tau\tau}$. For our analysis, we consider a benchmark value $\eta_{\alpha\alpha}=0.1~(\alpha=e,\mu,\tau)$ introducing one parameter at a time. We checked that these values lies within the bounds allowed by existing constraints on SNSI interaction Babu:2019iml . In fig. 1, we show the evolution of the effective mixing matrix elements involving $\nu_{e}$ ($U^{\rm eff}_{e1}$, $U^{\rm eff}_{e2}$, $U^{\rm eff}_{e3}$) as a function of matter density relevant for the supernova environment for neutrino energy $E=10$ MeV. The three columns show the presence of the SNSI parameters $\eta_{ee}$, $\eta_{\mu\mu}$ and $\eta_{\tau\tau}$ one at a time. The top row shows the results for NMO, whereas the bottom row corresponds to the IMO. In the presence of the SNSI parameter $\eta_{ee}$, we find that the effective mixing remains the same as in the standard case inside the supernova core ($\rho>10^{9}$ g/cc), i.e., $|U^{\rm eff}_{e3}|^{2}\simeq 1$, $|U^{\rm eff}_{e2}|^{2}\simeq|U^{\rm eff}_{e1}|^{2}\simeq 0$. However, in the presence of SNSI parameters $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$, we observe sharp changes in the effective mixing parameters compared to the standard case. For the NMO case, $|U^{\rm eff}_{e2}|\simeq 1$ and $|U^{\rm eff}_{e3}|\simeq 0$ at the supernova core, and $\nu_{e}$ aligns with effective $\nu_{2,m}$ states instead of $\nu_{3,m}$. Similarly, for the IMO case, $|U^{\rm eff}_{e1}|\simeq 1$, $|U^{\rm eff}_{e2}|\simeq 0$ at the core, and $\nu_{e}\simeq\nu_{1,m}$. We found that this happens due to the presence of a new crossing between the $\nu_{3,m}$ and $\nu_{2,m}$ ($\nu_{1,m}$) states at some value of the supernova matter density in the case of NMO (IMO). The resulting $\nu_{e}$ flux at the surface, $\Phi_{\nu_{e}}$ at the supernova envelope is written as
@@
$\Phi(\nu_{e})$ $=|U_{e1}|^{2}\Phi_{c}(\nu_{e})+(1-|U_{e1}|^{2})\,\Phi_{c}(\nu_{x})\,\,\,\,\,\,...\,\,{\rm for\ IMO}$ (14)
Note that in fig. 1, we use the value of SNSI parameters to be 0.1, however, our observations are valid for even smaller values of the SNSI parameters ($\eta_{\alpha\alpha}\approx 10^{-2}$).
@@ 4 Supernova neutrino events in next-generation neutrino experiments
Figure 2: The expected supernova events ($\nu_{e}$) distribution at DUNE as a function time after the core bounce. For the $\nu_{e}$ flux, we consider a supernova explosion with $25_{\odot}$ progenitor mass at a distance 10 kpc from the Earth. The blue line correspond to the case with standard CC interaction (SI) of neutrinos with background matter in supernova, whereas red line shows the same in the presence of additional SNSI interaction with $\eta_{\mu\mu}=0.1$ or $\eta_{\tau\tau}=0.1$. In the left (right) panel, event distribution are plotted with normal (inverted) mass ordering of neutrino. The values of the other oscillation parameters are taken from ref. Esteban:2024eli .
@@
In fig. 2, we show the expected $\nu_{e}$ event distribution at DUNE under both the standard interaction (SI) scenario and in the presence of SNSI characterized by the parameters $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ with a value of 0.1. Results are shown for both mass-ordering scenarios, illustrated in the two panels. In both NMO and IMO cases, the inclusion of SNSI parameters $\eta_{\mu\mu}=0.1$ or $\eta_{\tau\tau}=0.1$ significantly modifies the $\nu_{e}$ event distribution compared to the SI case. This deviation arises primarily because the effective mass eigenstate with which $\nu_{e}$ is aligned in the supernova core differs between the SI and SNSI cases. For the NMO case, $\nu_{e}\simeq\nu_{3,m}$ in the supernova core in the SI case, leading to an observed flux at Earth proportional to $|U_{e3}|^{2}$, considering $\nu_{e}$ flux dominates over the flux of other two flavors. In contrast, in the presence of SNSI, $\nu_{e}\simeq\nu_{2,m}$ at the supernova core, resulting in a $\nu_{e}$ flux at Earth proportional to $|U_{e2}|^{2}$. Since $|U_{e2}|^{2}>>|U_{e3}|^{2}$, we observe a greater number of events in the initial time after the core bounce. At later times, the contributions from the $\nu_{\mu}$ and $\nu_{\tau}$ become relevant and reduce the distinction between the two cases. In the IMO case, $\nu_{e}\simeq\nu_{2,m}$ and $\nu_{e}\simeq\nu_{1,m}$ at the supernova core in the standard interaction case and the presence of new interaction, respectively. Since $|U_{e1}|^{2}>|U_{e2}|^{2}$, we observe more events in the presence of the SNSI interaction than in the SI case.
@@
Figure 3: The expected supernova events ($\bar{\nu}_{e}$) distribution at Hyper-K with two identical detectors as a function time after the core bounce. For the $\bar{\nu}_{e}$ flux, we consider a supernova explosion with $25_{\odot}$ progenitor mass at a distance 10 kpc from the Earth. The blue line correspond to the case with standard CC interaction (SI) of neutrinos with background matter in supernova, whereas red line shows the same in the presence of additional SNSI interaction with $\eta_{\mu\mu}=0.1$ or $\eta_{\tau\tau}=0.1$. In the left (right) panel, event distribution are plotted with normal (inverted) mass ordering of neutrino. The values of the other oscillation parameters are taken from ref. Esteban:2024eli .
@@
One interesting observation is that the event distribution is the same for the SNSI case with NMO and the SI case with IMO. In both cases, the electron neutrino aligns with the same effective mass eigenstate, $\nu_{e}\simeq\nu_{2,m}$ at the source. Consequently, in the presence of new interaction for $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$, DUNE alone can not distinguish the two neutrino mass ordering scenario, as the SNSI introduces a new degeneracy in the mass ordering measurement. This is also true for the smaller values of $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ ($\simeq 10^{-2}$) since the high density of the supernova core gives the same effective mass eigenstate alignment as with the parameter equal to 0.1. The degeneracy that may occur at DUNE can be broken by including complementary observations of $\bar{\nu}_{e}$ events from a supernova explosion. As discussed earlier, the upcoming Hyper-Kamiokande (Hyper-K) experiment will be capable of detecting $\bar{\nu}_{e}$ via the inverse beta decay process. In Fig. 3, we show the expected $\bar{\nu}_{e}$ events in the Hyper-K detector. We observe a difference of around 10 events in each time bin for the NMO and IMO scenarios in the SI case, which arises due to the fact $\bar{\nu}_{e}$ aligns with $\bar{\nu}_{1,m}$ and $\bar{\nu}_{3,m}$ states, respectively for NMO and IMO. The event distributions remain the same in the presence of the scalar-mediated interaction parameters. It is because, unlike for neutrinos, no new level crossing happens for $\bar{\nu}_{e}$ flavor evolution in the presence of new physics parameter $\eta_{\mu\mu}$ and $\eta_{\tau\tau}$. However, unlike the DUNE, the distribution for the SNSI case with NMO is different from the distribution in the SI case with IMO. Therefore, a combined analysis of the event distribution from DUNE and Hyper-K could help distinguish between these two scenarios. If the $\nu_{e}$ event distribution observed at DUNE from a future supernova resembles the distribution shown by the red (blue) curve in the left (right) panel of Fig. 2, it would infer either NMO with SNSI or the IMO with SI. However, a complementary analysis of the $\bar{\nu}_{e}$ event distribution at Hyper-K may help resolve this degeneracy and identify the true underlying scenario. Other neutrino experiments, like JUNO, that measure $\bar{\nu}_{e}$ with an energy threshold similar to DUNE or Hyper-K may also help infer the true scenario.

### v2

@@ 3 Neutrino flavor evolution in core-collapse supernova in the presence of flavor conserving SNSI
We now investigate the impact of scalar-mediated non-standard interactions on neutrino flavor evolution in the supernova environment. Specifically, we focus on flavor-diagonal SNSI parameters: $\eta_{ee}$, $\eta_{\mu\mu}$, and $\eta_{\tau\tau}$. For our analysis, we consider two benchmark values $\eta_{\alpha\alpha}=0.1$ and $0.01~(\alpha=e,\mu,\tau)$ introducing one parameter at a time. We checked that these values lies within the bounds allowed by existing constraints on SNSI interaction Babu:2019iml . We also checked that neutrino propagation is adiabatic inside the supernova even in the presence of SNSI (see appendix A). In fig. 1, we show the evolution of the effective mixing matrix elements involving $\nu_{e}$ ($U^{\rm eff}_{e1}$, $U^{\rm eff}_{e2}$, $U^{\rm eff}_{e3}$) as a function of matter density relevant for the supernova environment for neutrino energy $E=10$ MeV for NMO. The three columns show the presence of the SNSI parameters $\eta_{ee}$, $\eta_{\mu\mu}$ and $\eta_{\tau\tau}$ one at a time. The top row shows results with SNSI strength $\eta_{\alpha\alpha}=0.1$, while the bottom row corresponds to the results with $\eta_{\alpha\alpha}=0.01$. In the presence of the SNSI parameter $\eta_{ee}$, we find that the effective mixing remains the same as in the standard case inside the supernova core where $\rho\gtrsim 10^{11}$ g/cc, i.e., $|U^{\rm eff}_{e3}|^{2}\simeq 1$, $|U^{\rm eff}_{e2}|^{2}\simeq|U^{\rm eff}_{e1}|^{2}\simeq 0$. However, in the presence of SNSI parameters $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$, we observe sharp changes in the effective mixing parameters compared to the standard case. For $\eta_{\mu\mu}=0.1$ or $\eta_{\tau\tau}=0.1$, $|U^{\rm eff}_{e2}|\simeq 1$ and $|U^{\rm eff}_{e1}|=|U^{\rm eff}_{e3}|\simeq 0$ at the supernova core ($\rho\gtrsim 10^{11}$ g/cc). This happens because of the sharp transition between the effective mixings $|U^{\rm eff}_{e3}|$ and $|U^{\rm eff}_{e2}|$ as seen in fig. 1  [^1] In fig. 1, $|U^{\rm eff}_{e2}|$ increases from 0 to 1 (while $|U^{\rm eff}_{e3}|$ decreases from 1 to 0) at a density $\rho=\rho_{\rm res}\simeq 4\times 10^{8}$ g/cc ($4\times 10^{10}$ g/cc) in the presence of $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ with strength 0.1 (0.01). This leads to a transition between $\nu_{3,m}$ and $\nu_{2,m}$ and results in $\nu_{e}\simeq\nu_{3,m}$ to $\nu_{e}\simeq\nu_{2,m}$ at $\rho>\rho_{\rm res}$., where the transition occurs for $\rho<10^{11}$ g/cc. As long as the sharp transition occurs outside of the core, essentially all of the $\nu_{e}$ flux produced in the neutronization phase is in the $\nu_{2,m}$ state. Now, at the supernova core, $\nu_{e}$ aligns with the $\nu_{2,m}$ ($\nu_{e}\simeq\nu_{2,m}$). In fig. 2, we show the same as in fig. 1, but for the inverted mass ordering of neutrinos. Similar to the NMO case, we observe a flip between the $|U^{\rm eff}_{e1}|$ and $|U^{\rm eff}_{e2}|$ in presence of SNSI parameters $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$. Now, $\nu_{e}$ aligns with the $\nu_{1,m}$ ($\nu_{e}\simeq\nu_{1,m}$) state inside the supernova core.
In figs. 1 and 2, we took the evolution of the effective mixing given the matter density inside the supernova, considering the strength of SNSI parameters as low as 0.01. As we go toward lower values of the SNSI parameters, the transition $\nu_{e}\simeq\nu_{2,m}$ ($\nu_{e}\simeq\nu_{1,m}$) from $\nu_{e}\simeq\nu_{3,m}$ ($\nu_{e}\simeq\nu_{2,m}$) states for NMO (IMO), which we call the “SNSI resonance”, occurs at a higher density. When the SNSI parameter $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ is too small, the resonance density is above the critical density $\rho_{c}=10^{14}$ g/cc, so resonance due to SNSI does not occur within the supernova. As noted above, we also required that the SNSI resonance occur outside of the core so that all $\nu_{e}$ in the neutronization phase align with a singe mass state $\nu_{i,m}$.
To estimate the minimum allowed value of the SNSI parameter for which the new resonance occurs within the supernova matter profile, we derive an expression for the resonance condition using a two-flavor approximation. Considering the $(e,\mu)$ pair in the Hamiltonian, and neglecting the elements from the vacuum Hamiltonian [^2] We checked that the value of elements of the vacuum Hamiltonian are several orders smaller than the interaction part of the Hamiltonian at the relevant matter densities.
[…]
In the above equation, $M_{\alpha\beta}~(\delta M_{\alpha\alpha})$ is the element of the vacuum Hamiltonian as defined in eq. (3) (eq. (7)). Diagonalizing the Hamiltonian, we derive the resonance condition
$2EV_{\rm CC}=(\delta M_{\mu\mu})^{2}\,.$ (14)
The above resonance condition for neutrinos gives the expression for the density at resonance,
$\rho_{\rm res}=\frac{2\sqrt{2}G_{F}Y_{e}E}{m_{n}|\Delta m^{2}_{31}|\eta^{2}_{\mu\mu}}\rho^{2}_{s}\,$, (15)
where $m_{n}$ is nucleon mass, and $Y_{e}$ is the electron fraction. As mentioned earlier, we take $\rho_{s}=10^{5}$ g/cc in this work. The resonance density depends on $\eta_{\mu\mu}^{2}$, so it is independent of the sign of $\eta_{\mu\mu}$. Since the new resonance density also depends on the absolute value of the mass-squared difference, we observe the new resonance for both the mass orderings of neutrinos. For NMO, $m_{2}$ and $m_{3}$ are the closest mass states, so we see a transition between $|U^{\rm eff}_{e2}|$ and $|U^{\rm eff}_{e3}|$. However, for IMO, $m_{2}$ and $m_{1}$ are the closest mass states, and we see a transition between $|U^{\rm eff}_{e1}|$ and $|U^{\rm eff}_{e2}|$. For antineutrinos, $V_{\rm CC}$ changes sign and the resonance condition in eq. (14) is not valid anymore, even for negative SNSI parameters, as it appears in the quadratic form in eq. (15). As a result, we do not observe any new resonance due to the SNSI in the case of antineutrinos.
To check the validity of the two-flavor approximation to derive eq. (15), we plot $\rho_{\rm res}$ as a function of $\eta_{\mu\mu}$ in the left panel of fig. 3 using $E=10$ MeV. The blue and red dots correspond to $\eta_{\mu\mu}=0.1$ and $0.01$, for which resonance densities are $\rho_{\rm res}\simeq 4.2\times 10^{8}$ g/cc and $4.2\times 10^{10}$ g/cc. In the right panel, we show the evolution of $|U^{\rm eff}_{e2}|$ at these two densities and at same neutrino energy, as indicated by the blue and red lines. We observe that the transition occurs at $\eta_{\mu\mu}\simeq 0.1$ and $\eta_{\mu\mu}\simeq 0.01$, respectively, which correspond to the blue and red dots in the left panel. There is a small deviation from the $\eta_{\mu\mu}$ equal to $0.1$ and $0.01$ due to the two-flavor approximation.
To have all the electron neutrinos produced during the neutronization burst phase in the new effective mass states ($\nu_{2,m}$ ($\nu_{1,m}$) state for NMO (IMO)), the SNSI resonance needs to occur outside the core of the supernova inside of which the neutronization phase occurs. Considering $\rho\simeq 10^{11}$ g/cc at the boundary of supernova core Jana:2024lfm , the dashed magenta line shows the minimum value of $\eta_{\mu\mu}$ ($\approx 0.006$) for which SNSI resonance for a 10 MeV neutrino occurs outside the supernova core. This means that for $\eta_{\mu\mu}\gtrsim 0.006$, the entire $\nu_{e}$ flux produced during the neutronization phase aligns with the $\nu_{2,m}$ ($\nu_{1,m}$) for NMO (IMO). We checked that for $E=40$ MeV, the maximum neutrino energy we consider in this work, this limit becomes $\eta_{\mu\mu}\gtrsim 0.01$. However, neutrino flux at such high energies is so small that the the number of events from these energies is negligible. The gray shaded region in the left panel shows the possible densities inside the supernova ($\rho~<~10^{14}$ g/cc). The minimum value of $\eta_{\mu\mu}$ that produces the new resonance inside the supernova is $\eta_{\mu\mu}({\rm min})\simeq 2\times 10^{-4}$. We checked that our two-flavor approximation of $\eta_{\mu\mu}({\rm min})$ matches the exact three-flavor scenario, as shown by the orange line in the right panel.
Figure 3: Left: For neutrinos, the resonance density $\rho_{\rm res}$ (eq. (15)) as a function of the SNSI strength $\eta_{\mu\mu}$ estimated analytically with two flavor approximation. The shaded region shows the possible matter densities inside the supernova ($\rho<10^{14}$ g/cc). The orange vertical line corresponds to the minimum value of $\eta_{\mu\mu}$, which keeps the new resonance inside the supernova. The magenta vertical line depicts the $\eta_{\mu\mu}$ for which resonance occurs at $\rho\simeq 10^{11}$ g/$c$, the density at the core boundary. The red and blue dots show the resonance densities that correspond to the $\eta_{\mu\mu}=0.01$ and $\eta_{\mu\mu}=0.1$, respectively. Right: Effective mixing $|U^{\rm eff}_{e2}|$ as function of the $\eta_{\mu\mu}$ for four benchmark value of the matter density inside the supernova: $\rho=4.2\times 10^{8}$ g/cc ($\rho_{res}$ for $\eta_{\mu\mu}=0.1$), $\rho=4.2\times 10^{10}$ g/cc ($\rho_{res}$ for $\eta_{\mu\mu}=0.01$), $\rho=4.2\times 10^{11}$ g/cc ($\rho_{res}$ for $\eta_{\mu\mu}=0.006$), and $\rho=10^{14}$ g/cc ($\rho_{res}$ for $\eta_{\mu\mu}=2\times 10^{-4}$ g/$c$). $|U^{\rm eff}_{e2}|$ is computed numerically with full three-flavor formalism. We consider neutrino energy $E=10$ MeV in both the panels.
The resulting $\nu_{e}$ flux at the surface, $\Phi_{\nu_{e}}$ at the supernova envelope is written as
@@
$\Phi(\nu_{e})$ $=|U_{e1}|^{2}\Phi_{c}(\nu_{e})+(1-|U_{e1}|^{2})\,\Phi_{c}(\nu_{x})\,\,\,\,\,\,...\,\,{\rm for\ IMO}\,.$ (17)
@@ 4 Supernova neutrino events in next-generation neutrino experiments
Figure 4: The expected supernova neutrino events ($\nu_{e}$) distribution at DUNE as a function time after the core bounce. The left (right) panels shows the distributions for NMO (IMO). In the top (bottom) panels, we use 10 kpc (20 kpc) as distance of the supernova from the Earth. For the $\nu_{e}$ flux, we consider a supernova explosion with $25_{\odot}$ progenitor mass at a distance 10 kpc from the Earth. The blue line correspond to the case with standard CC interaction (SI) of neutrinos with background matter in supernova, whereas red line shows the same in the presence of additional SNSI interaction with $\eta_{\mu\mu}=0.1$ or $\eta_{\tau\tau}=0.1$. The error bar shows the event rates considering flux and statistical uncertainties. The values of the other oscillation parameters are taken from ref. Esteban:2024eli .
@@
In fig. 4, we show the expected $\nu_{e}$ event distributions as a function of time after the core bounce at DUNE assuming either NMO (left panels) or IMO (right panels). We use two benchmark values of the distance of the supernova from Earth, 10 kpc and 20 kpc as shown in the top and bottom panels. We plot the distributions for the SI case and also in the presence SNSI parameters $\eta_{\mu\mu}$ or $\eta_{\tau\tau}$ with strength 0.1. For the NMO case, $\nu_{e}\simeq\nu_{3,m}$ in the supernova core in the SI case, leading to an observed flux at Earth proportional to $|U_{e3}|^{2}$, since the $\nu_{e}$ flux dominates over the flux of other two flavors. In contrast, in the presence of SNSI, $\nu_{e}\simeq\nu_{2,m}$ at the supernova core, resulting in a $\nu_{e}$ flux at Earth proportional to $|U_{e2}|^{2}$. Since $|U_{e2}|^{2}\gg|U_{e3}|^{2}$, we observe a greater number of events in the initial time after the core bounce. At later times, the contributions from the $\nu_{\mu}$ and $\nu_{\tau}$ become relevant and reduce the distinction between the two cases. In the IMO case, $\nu_{e}\simeq\nu_{2,m}$ and $\nu_{e}\simeq\nu_{1,m}$ at the supernova core in the SI case and in the presence of new interaction, respectively. Since $|U_{e1}|^{2}>|U_{e2}|^{2}$, we observe more events in the presence of the SNSI interaction than in the SI case.
Increasing the distance between the supernova and the Earth $R$ by a factor of two reduces the event rate in each time bin by a factor of $1/4$, as shown in the bottom panels, since the neutrino flux $f_{\nu_{\alpha},\oplus}$ at Earth is proportional to the $1/R^{2}$. We observe that for $R=20$ kpc, although the event rate reduces compared to $R=10$ kpc, DUNE can still differentiate between the SI and SNSI scenario in the initial time after the core bounce ($t\lesssim 15$ ms), considering the flux and statistical uncertainties, if the neutrino mass ordering is known. This difference is not prominent anymore for $R>20$ kpc, as event rates will be extremely small due to the $1/R^{2}$ suppression. Note that although we show our result for $\eta_{\mu\mu}=0.1$ or $\eta_{\tau\tau}=0.1$, our results are valid for even smaller value of the SNSI parameters, as long as the new resonance occurs outside the core of the supernova, which correspond to $\eta_{\mu\mu}\gtrsim 0.006$ or $\eta_{\tau\tau}\gtrsim 0.006$ (see fig. 3 and related discussion).
@@
Figure 5: The expected supernova neutrino events ($\bar{\nu}_{e}$) distribution at Hyper-K with two identical detectors as a function time after the core bounce. The left (right) panels shows the distributions for NMO (IMO). In the top (bottom) panels, we use $R=10$ kpc (20 kpc) as the distance of the supernova from the Earth. For the $\bar{\nu}_{e}$ flux, we consider a supernova explosion with $25_{\odot}$ progenitor mass. The blue line corresponds to the case with standard CC interaction (SI) of neutrinos with background matter in supernova, whereas the overlapping red dashed line shows the same in the presence of additional SNSI interaction with $\eta_{\mu\mu}=0.1$ or $\eta_{\tau\tau}=0.1$. The error bar shows the event rates considering flux and statistical uncertainties. The values of the other oscillation parameters are taken from ref. Esteban:2024eli .
@@
One interesting observation is that the event distribution is the same for the SNSI case with NMO and the SI case with IMO. In both cases, the electron neutrino aligns with the same effective mass eigenstate, $\nu_{e}\simeq\nu_{2,m}$ at the source. Consequently, the presence of a new interaction ($\eta_{\mu\mu}$ or $\eta_{\tau\tau}$) can brings in a new degeneracy in mass ordering measurement, and DUNE alone can not distinguish the two scenarios. The degeneracy that may occur at DUNE can be broken by including complementary observations of $\bar{\nu}_{e}$ events from a supernova explosion. As discussed earlier, the upcoming Hyper-Kamiokande (Hyper-K) experiment will be capable of detecting $\bar{\nu}_{e}$ via the inverse beta decay process if the $\bar{\nu}_{e}$ flux is large enough. In Fig. 5, we show the expected $\bar{\nu}_{e}$ events in the Hyper-K detector for the $25M_{\odot}$ supernova model. For $R=10$ kpc, we observe a difference of around 10 events in each time bin for the NMO and IMO scenarios in the SI case, which arises because $\bar{\nu}_{e}$ aligns with $\bar{\nu}_{1,m}$ and $\bar{\nu}_{3,m}$ states, respectively for NMO and IMO. For $R=20$ kpc, the event rate is reduced by a factor of $1/4$ as expected, and the difference between event distribution for two mass ordering scenarios almost vanishes once we consider the flux and statistical uncertainties. The event distributions remain the same in the presence of the scalar-mediated interaction parameters. It is because, unlike for neutrinos, no new level crossing happens for $\bar{\nu}_{e}$ flavor evolution in the presence of new physics parameter $\eta_{\mu\mu}$ and $\eta_{\tau\tau}$. However, unlike the DUNE, the distribution for the SNSI case with NMO is different from the distribution in the SI case with IMO. Therefore, a combined analysis of the event distribution from DUNE and Hyper-K could help distinguish between these two scenarios as long as $R\lesssim 10$ kpc. If the $\nu_{e}$ event distribution observed at DUNE from a future supernova resembles the distribution shown by the red (blue) curve in the left (right) panel of Fig. 4, it would infer either NMO with SNSI or the IMO with SI. However, for $R\lesssim 10$ kpc, a complementary analysis of the $\bar{\nu}_{e}$ event distribution at Hyper-K may help resolve this degeneracy and identify the true underlying scenario. Other neutrino experiments, like JUNO, that measure $\bar{\nu}_{e}$ with an energy threshold similar to DUNE or Hyper-K may also help infer the true scenario.